Determine the total charge entering a terminal between t = 1 and t = 2 s and t = 2 s if the current passing the terminal is I = (3t2 -t) A.



1
Expert's answer
2021-09-06T11:34:57-0400

"dq=Idt,"

"q=\intop_1^2(3t^2-t)dt=(t^3-0.5t^2)|_1^2=5.5~C."